OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5)

OpenStudy (gorv):

fog(-5)=f(g(-5)) g(-5)=-4*(-5)+3 =20+3=23

OpenStudy (gorv):

f(23)=-2*23-7=-46-7=-53

OpenStudy (anonymous):

@gorv Let f(x) = x2 + 6 and g(x) = x+8/x Find ( g o f)(-7)